Jessica is older than Janet. Their sum of their ages is 38 years. Twice janets age plus jessicas age is 51. How old is janet?
I tried:
Let x be greater
x=Jessica
y = Janet
x+ y = 38
2y+x=51



Answer by jim_thompson5910(35256)

You have the correct equations. Now you just need to solve.


x+ y = 38

y = 38-x
-------------------------------------------------------
2y+x=51

2(38-x)+x=51


Here you now have an equation with one variable. I'll let you solve it.

Once you find the value of x, use it with y = 38-x to find the value of y.